OpenStudy (anonymous):

What is the value of c that satisfies the Mean Value Theorem (MVT) for f(x)=x^3-3x^2+x on [0, 3]?

OpenStudy (dumbcow):

f'(c) = 1/3[f(3) -f(0)] f'(c) = 1 **f'(x) = 3x^2-6x+1 ** 3c^2 -6c +1 = 1 3c^2 = 6c c^2 = 2c c = 0,2, since 0 is one of the endpoints, then c=2